OpenStudy (anonymous):

Simplify: (-3t – 5)8

OpenStudy (lalaly):

-24t-40

OpenStudy (anonymous):

-24t - 40

OpenStudy (anonymous):

\[-3t\times 8-5\times 8=-24t-40\]

OpenStudy (anonymous):

sattelite73 answer my question

OpenStudy (anonymous):

please?

OpenStudy (anonymous):

and don't forget to ask your math teacher the question "in what sense is \[-24t-40\] "simpler" than \[(-3t-5)8\] ? because the answer is it is NOT